Hours/month Total Utility Windsurfing Total Utility Snorkeling
0_____________0____________________0
1_____________60___________________20
2_____________110__________________38
3_____________150__________________53
4_____________180__________________64
5_____________200__________________70

Devon's total income rises from $35 to $55, ceteris paribus. Given that Devon changes his of choice of windsurfing and snorkeling, what is his income elasticity for windsurfing?

The correct answer is 0.64. However, I do not know the steps to find this value.
1
Expert's answer
2012-07-24T11:58:44-0400
IEd = (Y1+Y2)/(Q1+Q2) * (Q2 - Q1)/(Y2 - Y1) = 90/(Q1+Q2) * (Q2 - Q1)/20 = 4.5*(Q2 - Q1)/(Q2 + Q1) = 4.5*(4 - 3)/(4+3) = 0.64
As the total utility for windsurfing is higher, than for snorkeling, Devon will increase spending time only for windsurfing.